# Towards Transferable Targeted 3D Adversarial Attack in the Physical World

### v1

@@ Abstract
Compared with transferable untargeted attacks, transferable targeted adversarial attacks could specify the misclassification categories of adversarial samples, posing a greater threat to security-critical tasks. In the meanwhile, 3D adversarial samples, due to their potential of multi-view robustness, can more comprehensively identify weaknesses in existing deep learning systems, possessing great application value. However, the field of transferable targeted 3D adversarial attacks remains vacant. The goal of this work is to develop a more effective technique that could generate transferable targeted 3D adversarial examples, filling the gap in this field. To achieve this goal, we design a novel framework named TT3D that could rapidly reconstruct from few multi-view images into Transferable Targeted 3D textured meshes. While existing mesh-based texture optimization methods compute gradients in the high-dimensional mesh space and easily fall into local optima, leading to unsatisfactory transferability and distinct distortions, TT3D innovatively performs dual optimization towards both feature grid and Multi-layer Perceptron (MLP) parameters in the grid-based NeRF space, which significantly enhances black-box transferability while enjoying naturalness. Experimental results show that TT3D not only exhibits superior cross-model transferability but also maintains considerable adaptability across different renders and vision tasks. More importantly, we produce 3D adversarial examples with 3D printing techniques in the real world and verify their robust performance under various scenarios.
@@ 1 Introduction
Despite the unprecedented performance of deep neural networks (DNNs) in numerous tasks, including image classification and object detection, these models are vulnerable to adversarial examples [33, 9, 21, 19, 4, 46, 37, 41, 42, 5, 44, 27, 43]. These adversarial examples are shown to have good transferability across models [16, 4, 46], emerging as a key concern since they can evade black-box models in practical applications. The majority of transferable attacks are untargeted, aiming to induce misclassification of the target model. However, transferable targeted attacks, which mislead DNNs to produce predetermined misclassifications, can lead to a more severe threat in real-world applications. Crafting such transferable targeted attacks is particularly challenging because they not only need to achieve a specific misclassification but also must avoid overfitting to ensure effective attacks across various models, which requires further investigation.
@@
Target Task Classify Classify Classify Classify Detect Detect $*{}^{*}$ $*{}^{*}$$*{}^{*}$
@@
Targeted
Transferability
Naturalness
Physical Attack
Table 1: A comparison among different 3D attack methods regarding target vision tasks, used data, 3D attack types, transferability, whether conducting targeted attacks, naturalness, and whether conducting attacks in the physical world. $*{}^{*}$ represents the transferred task .
@@
Figure 2: An overview of our TT3D framework. We first utilize 3D multi-view reconstruction technology, i.e., grid-based NeRF with marching cubes techniques to obtain the initial clean 3D mesh. Then, we perform adversarial fine-tuning in the textual parameter space of grid-based NeRF instead of directly altering the texture $\mathcal{T}$, supplemented with geometric perturbations at vertex positions $\mathcal{V}$. To ensure the naturalness simultaneously, we add constrains to the distance between the 3D adversarial samples and the initial ones in terms of both texture and geometric structure when performing optimization .
@@
  • Experimental results demonstrate that our 3D adversarial object can easily be misclassified as a given label across various victim models, renders, and tasks. Moreover, we produce 3D adversarial objects in the real world by utilizing 3D printing technology and validate their robust performance under various settings, like different viewpoints and backgrounds.
@@ 3.2 Problem Formulation
$\mathop{\text{min}}\limits_{\mathcal{V}^{*},\mathcal{T}^{*}}\mathbb{E}_{\boldsymbol{\mathbf{v}}\in V}L_{f}(\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv}),y^{*})+\beta\cdot\mathcal{R}(\mathcal{M}_{adv},\mathcal{M}),$
where $\text{where }\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv})=\mathcal{S}(\mathcal{V}^{*},\mathcal{T}^{*},\mathcal{F},\boldsymbol{\mathbf{v}}),$ (3)
where $\hat{I}_{v}(\mathcal{M}_{adv})$ represents the rendered image of $\mathcal{M}_{adv}$ under the viewpoint $\boldsymbol{\mathbf{v}}$ and $V$ is the feasible distribution. $L_{f}$ is the cross-entropy loss [2] that facilitates the misclassification of $\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv})$ to $y^{*}$, $R$ is the regularization for minimizing a perceptibility distance between $\mathcal{M}_{adv}$ and $\mathcal{M}$ and $\beta$ is a balancing hyperparameter between these two losses.
To ensure the feasibility of our 3D adversarial example in the real world, we also introduce an EOT scheme (as detailed in Sec. 3.5) to help bridge the gap from “digital” to “physical”. However, the mesh-based optimization by following the objective function (3) needs to calculate gradients in high-dimensional mesh space due to thousands of points in each 3D object. It will easily distort the natural textural appearance of the textured mesh and is trapped into the overfitting with unsatisfactory transferability.
@@ 3.3 Dual Optimization in NeRF Parameter Space
To sum up, we target the grid-based NeRF’s inner parameters: $\left(\Theta_{G_{tex}},\Theta_{M_{tex}}\right)$ instead of vertex colors $\mathcal{T}$ in appearance, accompanying the optimization to mesh’s vertex coordinates, thus the objective function (3) becomes:
$\mathop{\text{min}}\limits_{\mathcal{V}^{*},\Theta_{G_{tex}}^{*},\Theta_{M_{tex}^{*}}}\mathbb{E}_{\boldsymbol{\mathbf{v}}\in V}L_{f}(\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv}),y^{*})+$
$\beta\cdot\mathcal{R}(\mathcal{M}_{adv},\mathcal{M}).$ (5)
@@ 3.4 Regularization for Naturalness
$R(\mathcal{M}_{adv},\mathcal{M})=\underbrace{\lambda_{1}R_{rgb}(\mathcal{M}_{adv},\mathcal{M})}_{\text{appearance}}+$
$\underbrace{\lambda_{2}R_{cd}(\mathcal{M}_{adv},\mathcal{M})+\lambda_{3}R_{lap}(\mathcal{M}_{adv})+\lambda_{4}R_{edge}(\mathcal{M}_{adv})}_{\text{geometry}},$ (7)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{4}$ are hyperparameters that represent weights of $R_{rgb}$, $R_{cd}$, $R_{lap}$, and $R_{edge}$, respectively. $R_{cd}$, $R_{lap}$, and ,$R_{edge}$ are detailed in Supplemental Materials.
@@ 3.5 Physical Attack
To realize physically feasible 3D adversarial examples, we need to ensure their robustness against complex transformations in the physical world, including 3D rotations, affine projections, color discrepancies, etc. A prevalent technique that we utilize here is the Expectation Over Transformation (EOT) algorithm [1], which optimizes the adversarial example across a distribution of varying transformations in both 2D and 3D spaces. Specifically, we effectively merge diverse 3D transformations, such as pose, distance, and viewpoint shifts, during rendering with 2D transformations such as contrast and blurring on rendered images. With EOT, the rendered images for optimization in Eq. 5 can be as below:
@@
MLP+Grid 70.98 51.61 59.59 57.59 50.25 43.81 $96.74^{*}$ 40.59 45.06 55.15 36.49
Table 2: The attack success rates(%) of 3d adversarial examples generated by different methods including the enhanced mesh-based, the mlp-only, the grid-only, and our dual optimization method under random viewpoints against ResNet-50 (RN-50), ResNet-101 (RN-101), ResNet-152 (RN-152), VGG-16, VGG-19, Inception-v3 (Inc-v3), DenseNet-121 (DN-121), EfficientNet-B0 (EN-B0), MobileNet-v2 (MN-v2), Swin-B, and VIT-B/16. The adversarial examples are learned against the surrogate models ResNet-101 and DenseNet-121.
@@ 4.1 Experiment Settings
Victim models. Two typical classifier models, ResNet-101 [10] and DenseNet-121 [11] are chosen as surrogate models to perform attacks. Other models to test transferability consist of the CNN-based ResNet-50 [10], ResNet-152 [10], VGG-16 [29], VGG19 [29], Inceprion-v3 [34], EfficientNet-B0 [35], MobileNet-V2 [28]and the Transformer-based Swin-B [17], ViT-B/16 [7]. For testing the cross-render transferability, we choose two commercial rendering software: MeshLab and Blender. For testing the cross-task transferability, we choose two other tasks: zero-shot detection [15] and image caption [13] due to their non-restricted to types of objects.
@@
TT3D hyperparameters. $\beta$ in Eq. 5 is $10^{3}$, the epochs are 250, $\lambda_{1}$ is $1$, $\lambda_{2}$ is $3000$, $\lambda_{3}$ is $10^{-3}$, $\lambda_{4}$ is $10^{-2}$.
@@ 4.2.1 Basic Results
Source Model Render Victim Model
@@
Table 3: The attack success rates(%) of 3d adversarial examples with different renders against ResNet-50 (RN-50), ResNet-101 (RN-101), ResNet-152 (RN-152), VGG-16, VGG-19, Inception-v3 (Inc-v3), DenseNet-121 (DN-121), EfficientNet-B0 (EN-B0), MobileNet-v2 (MN-v2), Swin-B, and VIT-B/16. The adversarial examples are learned against surrogate models ResNet-101 and DenseNet-121.
@@
Tex+Geo 70.98 51.61 59.59 57.59 50.25 43.81 $\bm{96.74^{*}}$ 40.59 45.06 55.15 36.49
Table 4: The attack success rates(%) of tex-only and tex+geo(sub) optimization methods against ResNet-50 (RN-50), ResNet-101 (RN-101), ResNet-152 (RN-152), VGG-16, VGG-19, Inception-v3 (Inc-v3), DenseNet-121 (DN-121), EfficientNet-B0 (EN-B0), MobileNet-v2 (MN-v2), Swin-B, and VIT-B/16. The adversarial examples are learned against the surrogate models ResNet-101 and DenseNet-121.
@@ 4.2.2 Cross-render Transferability
Figure 6: Visiual examples of printed 3D adversarial objects towards the target label under different backgrounds (B-1,B-2, B-3) and viewpoints in the physical world. The first row is learned against ResNet-101 and the second row is learned against DenseNet-121.
4.2.3 Cross-task Transferability
To verify the transferability of our adversarial examples across different tasks, we select zero-shot detection and image caption for testing due to their broad applicability across a wide range of object types.
Zero-shot detector. We utilize the state-of-the-art model of Liu et al. [15] for zero-shot detection. In our tests, we eliminate background distractions by using a plain white background, focusing solely on the detector’s ability to identify the target label in the rendered adversarial sample images. A detection threshold greater than 0.5 is considered successful. Our experimental results reveal that even in cross-task tests with a plain white background, our method achieved a success rate of 76.94%, as evidenced by some successful examples shown in Fig. 5.
@@ B-1 B-2 B-3
Table 5: The attack success rates(%) of printed adversarial meshes against the ResNet-101 and DenseNet-121 under different backgrounds: B-1, B-2, B-3 and various viewpoints in the real wolrd.

### v3

@@ Abstract
Compared with transferable untargeted attacks, transferable targeted adversarial attacks could specify the misclassification categories of adversarial samples, posing a greater threat to security-critical tasks. In the meanwhile, 3D adversarial samples, due to their potential of multi-view robustness, can more comprehensively identify weaknesses in existing deep learning systems, possessing great application value. However, the field of transferable targeted 3D adversarial attacks remains vacant. The goal of this work is to develop a more effective technique that could generate transferable targeted 3D adversarial examples, filling the gap in this field. To achieve this goal, we design a novel framework named TT3D that could rapidly reconstruct from few multi-view images into Transferable Targeted 3D textured meshes. While existing mesh-based texture optimization methods compute gradients in the high-dimensional mesh space and easily fall into local optima, leading to unsatisfactory transferability and distinct distortions, TT3D innovatively performs dual optimization towards both feature grid and Multi-layer Perceptron (MLP) parameters in the grid-based NeRF space, which significantly enhances black-box transferability while enjoying naturalness. Experimental results show that TT3D not only exhibits superior cross-model transferability but also maintains considerable adaptability across different renders and vision tasks. More importantly, we produce 3D adversarial examples with 3D printing techniques in the real world and verify their robust performance under various scenarios. Code is publicly available at: hhttps://github.com/Aries-iai/TT3D
@@ 1 Introduction
Despite the unprecedented performance of deep neural networks (DNNs) in numerous tasks, including image classification [27] and object detection [57, 13], these models are vulnerable to adversarial examples [35, 9, 22, 20, 4, 51, 39, 44, 45, 5, 48, 29, 46, 47, 49]. These adversarial examples are shown to have good transferability across models [17, 4, 51], emerging as a key concern since they can evade black-box models in practical applications. The majority of transferable attacks are untargeted, aiming to induce misclassification of the target model. However, transferable targeted attacks, which mislead DNNs to produce predetermined misclassifications, can lead to a more severe threat in real-world applications. Crafting such transferable targeted attacks is particularly challenging because they not only need to achieve a specific misclassification but also must avoid overfitting to ensure effective attacks across various models, which requires further investigation.
@@
Target Task Classify Classify Classify Classify Detect Detect Detect+Segment Classify+Detect+Caption
@@
[…]
Table 1: A comparison among different 3D attack methods regarding target vision tasks, used data, 3D attack types, transferability, whether conducting targeted attacks, naturalness, and whether conducting attacks in the physical world. represents the transferred task .
@@
Figure 2: An overview of our TT3D framework. We first utilize 3D multi-view reconstruction technology, i.e., grid-based NeRF with marching cubes techniques to obtain the initial clean 3D mesh. Then, we perform adversarial fine-tuning in the textual parameter space of grid-based NeRF instead of directly altering the texture $\mathcal{T}$, supplemented with geometric perturbations at vertex positions $\mathcal{V}$. To ensure the naturalness simultaneously, we add constraints to the distance between the 3D adversarial samples and the initial ones in terms of both texture and geometric structure when performing optimization .
@@
  • Experimental results reveal that our 3D adversarial object can easily be misclassified as a given label across various victim models, renders, and tasks. Moreover, we produce 3D adversarial objects in the real world by utilizing 3D printing and validate their robust performance under various settings, like different viewpoints and backgrounds.
@@ 3.2 Problem Formulation
$\mathop{\text{min}}\limits_{\mathcal{V}^{*},\mathcal{T}^{*}}\mathbb{E}_{\boldsymbol{\mathbf{v}}\in V}L_{f}(\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv}),y^{*})+\beta\cdot\mathcal{R}(\mathcal{M}_{adv},\mathcal{M}),$ (3)
where $\text{where }\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv})=\mathcal{S}(\mathcal{V}^{*},\mathcal{T}^{*},\mathcal{F},\boldsymbol{\mathbf{v}}),$
where $\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv})$ represents the rendered image of $\mathcal{M}_{adv}$ under the viewpoint $\boldsymbol{\mathbf{v}}$ and $V$ is the feasible distribution. $L_{f}$ is the cross-entropy loss [2] that facilitates the misclassification of $\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv})$ to $y^{*}$, $R$ is the regularization for minimizing a perceptibility distance between $\mathcal{M}_{adv}$ and $\mathcal{M}$ and $\beta$ is a balancing hyperparameter between these two losses.
To ensure the feasibility of our 3D adversarial example in the real world, we also introduce an EOT scheme (as detailed in Sec. 3.5) to help bridge the gap from “digital" to “physical". However, the mesh-based optimization by following the objective function (3.2) needs to calculate gradients in high-dimensional mesh space due to thousands of points in each 3D object. It will easily distort the natural textural appearance of the textured mesh and is trapped into the overfitting with unsatisfactory transferability.
@@ 3.3 Dual Optimization in NeRF Parameter Space
To sum up, we target the grid-based NeRF’s inner parameters: $\left(\Theta_{G_{tex}},\Theta_{M_{tex}}\right)$ instead of vertex colors $\mathcal{T}$ in appearance, accompanying the optimization to mesh’s vertex coordinates, thus the objective function (3.2) becomes:
$\mathop{\text{min}}\limits_{\mathcal{V}^{*},\Theta_{G_{tex}}^{*},\Theta_{M_{tex}^{*}}}\mathbb{E}_{\boldsymbol{\mathbf{v}}\in V}L_{f}(\hat{I}_{\boldsymbol{\mathbf{v}}}(\mathcal{M}_{adv}),y^{*})+$ (5)
$\beta\cdot\mathcal{R}(\mathcal{M}_{adv},\mathcal{M}).$
@@ 3.4 Regularization for Naturalness
$R(\mathcal{M}_{adv},\mathcal{M})=\lambda_{1}R_{rgb}(\mathcal{M}_{adv},\mathcal{M})+$ (7)
$\lambda_{2}R_{cd}(\mathcal{M}_{adv},\mathcal{M})+\lambda_{3}R_{lap}(\mathcal{M}_{adv})+\lambda_{4}R_{edge}(\mathcal{M}_{adv}),$
where $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{4}$ are hyperparameters that represent weights of $R_{rgb}$, $R_{cd}$, $R_{lap}$, and $R_{edge}$, respectively.
@@ 3.5 Physical Attack
To realize physically feasible 3D adversarial examples, we need to ensure their robustness against complex transformations in the physical world, including 3D rotations, affine projections, color discrepancies, etc. A prevalent technique that we utilize here is the Expectation Over Transformation (EOT) algorithm [1], which optimizes the adversarial example across a distribution of varying transformations in both 2D and 3D spaces. Specifically, we effectively merge diverse 3D transformations, such as pose, distance, and viewpoint shifts, during rendering with 2D transformations such as contrast and blurring on rendered images. With EOT, the rendered images for optimization in Sec. 3.3 can be as below:
@@
[…]
Table 2: The ASR(%) of 3d adversarial examples generated by different methods including the enhanced mesh-based, the mlp-only, the grid-only, and our dual optimization method under random viewpoints against ResNet-50 (RN-50), ResNet-101 (RN-101), ResNet-152 (RN-152), VGG-16, VGG-19, Inception-v3 (Inc-v3), DenseNet-121 (DN-121), EfficientNet-B0 (EN-B0), MobileNet-v2 (MN-v2), Swin-B, and VIT-B/16. The adversarial examples are learned against the surrogate models ResNet-101 and DenseNet-121.
@@ 4.1 Experiment Settings
Victim models. Two typical classifier models, ResNet-101 [10] and DenseNet-121 [11] are chosen as surrogate models to perform attacks. Other models to test transferability consist of the CNN-based ResNet-50 [10], ResNet-152 [10], VGG-16 [31], VGG19 [31], Inceprion-v3 [36], EfficientNet-B0 [37], MobileNet-V2 [30]and the Transformer-based Swin-B [18], ViT-B/16 [7]. For testing the cross-render transferability, we choose two commercial rendering software: MeshLab and Blender. For testing the cross-task transferability, we choose two other tasks: zero-shot detection [16] and image caption [14] due to their non-limitation to types of objects.
@@
TT3D hyperparameters. $\beta$ in Sec. 3.3 is $10^{3}$, the epochs are 250, $\lambda_{1}$ is $1$, $\lambda_{2}$ is $3000$, $\lambda_{3}$ is $10^{-3}$, $\lambda_{4}$ is $10^{-2}$.
@@ 4.2.1 Basic Results
Source Model Methods Victim Model
@@
Table 3: The ASR(%) of 3d adversarial examples with different renders against RN-50, RN-101, RN-152, VGG-16, VGG-19, Inc-v3, DN-121, EN-B0, MN-v2, Swin-B, and VIT-B/16. The adversarial examples are learned against surrogate models RN-101 and DN-121.
@@
[…]
Table 4: The ASR(%) of tex-only and tex+geo(sub) optimization methods against RN-50, RN-101, RN-152, VGG-16, VGG-19, Inc-v3, DN-121, EN-B0, MN-v2, Swin-B, and VIT-B/16. The adversarial examples are learned against the surrogate models RN-101 and DN-121.
@@ 4.2.2 Cross-render Transferability
4.2.3 Cross-task Transferability
To verify the transferability of our adversarial examples across different tasks, we select zero-shot detection and image caption for testing due to their broad applicability.
Zero-shot detector. We utilize the state-of-the-art model of Liu et al. [16] for zero-shot detection. In our tests, we eliminate background distractions by using a plain white background, focusing solely on the detector’s ability to identify the target label in the rendered adversarial sample images. A detection threshold greater than 0.5 is considered successful. Experimental results reveal that even in this setting, our method achieved a success rate of 76.94%, as evidenced by some successful examples shown in Fig. 5.
@@ 4.2.3 Cross-task Transferability
Figure 6: Visual examples of printed 3D adversarial objects towards the target label under different backgrounds (B-1,B-2, B-3) and viewpoints in the physical world. The first row is learned against ResNet-101 and the second row is learned against DenseNet-121.
@@ B-1 B-2 B-3
Table 5: The ASR(%) of printed adversarial meshes against the RN-101 and DN-121 under different backgrounds: B-1, B-2, B-3 and various viewpoints in the real wolrd.
@@ 5 Conclusion
Acknowledgement
This work was supported in part by the Project of the National Natural Science Foundation of China (Nos. 62076018, U2341228, 62276149) and in part by the Fundamental Research Funds for the Central Universities.
Supplementary Material
Appendix A Physical Implementation Details
For the implementation of physical attacks, here we detail the specific process including the used equipment. The whole implementation process could be split into three stages: Preprocess—3D Print—Capture and Test.
Preprocess. Before 3D printing, there is still a need for some preprocessing of the generated adversarial 3D mesh. This is because we have parameterized the appearance information into the parameters of grid-based NeRF. Hence, it is necessary to convert this information back into the texture maps required for 3D printing. Specifically, to extract the appearance as texture images, we first unwrap the UV coordinates of $\mathcal{M}_{adv}$ using XAtlas [53]. Subsequently, we bake the surface’s color into an image of the texture map corresponding to the UV coordinates, which could be used for the following 3D printing.
3D Print. Then, we print the 3D adversarial examples generated by TT3D in the form of textured meshes using well-established 3D printers. In our case, we utilize the J850™ Digital Anatomy™ 3D Printer [^1] https://www.stratasys.com/en/3d-printers/printer-catalog/polyjet/j850-digital-anatomy/, a classic 3D printer known for its versatility and precision. This printer can accurately reproduce both the textures and geometries of our adversarial samples, offering a wide range of material choices and color options. Here, we print a total of 20 objects, including 10 against the surrogate model ResNet-101 and 10 against the surrogate DenseNet-121.
Capture and Test. In the last stage, we aim to validate the performance of the 3D adversarial samples generated by TT3D across different viewpoints and backgrounds in the physical world. Thus, as mentioned in the manuscript, we place the 3D adversarial object on the given surface (with different backgrounds B-1, B-2, and B-3) and slowly circle them with a smartphone for about $360^{\circ}$, excluding the bottom part. which lasts approximately 20 seconds per object in each setting, capturing 10 frames per second, resulting in a total of 200 frames. Finally, we could calculate the attack success rate as the final test result, which is determined by the proportion of successful frames.
Appendix B More Experiments
B.1 Choice for Epoch
To ensure convergence and prevent overfitting, selecting an appropriate epoch for TT3D is crucial. Thus, following the same settings as described in Section 4.1 of the manuscript, we have visualized the variation curve of the total loss value with respect to the epoch for the attack surrogate model ResNet-101 in Fig. 8. The figure reveals that when the epoch reaches the number of 250, the total loss value gradually stabilizes, signifying the achievement of convergence. To prevent overfitting, we have thus chosen 250 as the final epoch for optimization.
Figure 8: The total loss of TT3D with respect to the epoch against the surrogate model ResNet-101 with different training iterations.
B.2 Evaluation for Naturalness
[…]
Initial 0.9821 39.00 0.0293
Mesh-based 0.8259 13.79 0.1912
TT3D(RN-101) $\beta=10^{2}$ 0.8741 22.54 0.1447
$\beta=10^{3}$ 0.9039 27.96 0.1153
$\beta=10^{4}$ 0.9486 33.26 0.0874
Table 6: The results of quantitative naturalness metrics: PSNR, SSIM, and LPIPS scores for the initially reconstructed clean images, adversarial samples generated by the previous mesh-based method, and the corresponding samples produced by our TT3D method under $\beta=10^{2}$, $\beta=10^{3}$, and $\beta=10^{4}$. Higher SSIM and PSNR values indicate better performance, while a lower LPIPS score is preferable.
Figure 9: The attack success rate(%) of 3d adversarial examples generated by TT3D under $\beta=10^{2}$, $\beta=10^{3}$, and $\beta=10^{4}$ against ResNet-50 (RN-50), ResNet-101 (RN-101), ResNet-152 (RN-152), VGG-16, VGG-19, Inception-v3 (Inc-v3), DenseNet-121 (DN-121), EfficientNet-B0 (EN-B0), MobileNet-v2 (MN-v2), Swin-B, and VIT-B/16. The 3d adversarial examples are learned against surrogate models ResNet-101. Each bar chart is accompanied by a visual example in the lower right position, showcasing the appearance and geometry of a 3d adversarial sample (attack ‘bread’ to ‘stove’ ) generated by TT3D under the corresponding $\beta$ value.
Source Model Background Victim Model
[…]
DenseNet-121 B-1 39.25 34.70 36.35 29.70 26.55 32.15 $79.30^{*}$ 31.20 33.75 43.20 21.20
[…]
Table 7: The attack success rates(%) of 3d printed adversarial objects with different backgrounds in the physical world against ResNet-50 (RN-50), ResNet-101 (RN-101), ResNet-152 (RN-152), VGG-16, VGG-19, Inception-v3 (Inc-v3), DenseNet-121 (DN-121), EfficientNet-B0 (EN-B0), MobileNet-v2 (MN-v2), Swin-B, and VIT-B/16. The 3d adversarial objects are learned against the surrogate model ResNet-101 and DenseNet-121 and produced with 3d printing techniques for physical implementation.
To more objectively evaluate the naturalness of 3D adversarial samples generated by TT3D, three metrics for assessing image quality are employed here: Structural Similarity Index [42] (SSIM), Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity [55] (LPIPS). SSIM evaluates perceptual degradation by analyzing structural, luminance, and contrast changes. PSNR measures pixel-level accuracy between original and distorted images, with higher values indicating better fidelity. LPIPS, leveraging deep learning (vgg used here), evaluates perceptual similarity at the patch level, reflecting the human visual system’s response to image variations. Specifically, we compare the above three quantitative metrics of the initially reconstructed clean images, adversarial samples generated by the previous mesh-based method, and the corresponding ones produced by our TT3D method under various $\beta$ values. $\beta$ is the weight of the regularization for naturalness in TT3D. Experimental results are listed in Tab. 6, revealing that our TT3D significantly outperforms the mesh-based method even under different $\beta$ values and exhibits an acceptable decline compared to the clean images, thereby confirming the superior naturalness of TT3D.
B.3 Effects of $\beta$
As mentioned in the manuscript, $\beta$, an adjustable weight for the regularization $\mathcal{R}$ responsible for naturalness, has a significant impact on both naturalness and attack performance. To perform a more comprehensive analysis of $\beta$’s effects, here we measure the attack success rate of TT3D (against the surrogate model ResNet-101) across various black-box classifiers with different $\beta$ values of $10^{2}$, $10^{3}$, and $10^{4}$. Additionally, we provide a visual example of the adversarial samples generated under these settings. Experimental results, as presented in Fig. 9, where we can see that: 1) At $\bm{\beta=10^{2}}$, when $\beta$ is relatively low, the success rate across different models is predominantly above 80%. However, this comes at the cost of a certain degree of naturalness, both in appearance and geometry. 2) At $\bm{\beta=10^{4}}$, under a strong regularization weight, we still observe considerable success rates, with the lowest being above 30%. 3) At $\bm{\beta=10^{3}}$, there is a relative balance between success rate and naturalness, with both metrics demonstrating objectively favorable outcomes. Consequently, this $\beta$ value, i.e., $\beta=10^{3}$ is chosen as the final implementation parameter for TT3D.
Figure 10: The total of 100 different 3D adversarial objects, generated by TT3D to attack the surrogate models ResNet-101. The above image of each object is captured under a single, random viewpoint, and in the experiments, we capture 100 different viewpoints for calculating the ASR. The green text under each object is the clean label, and the red text is the target label, which is also randomly chosen from the labels in Imagenet. The above images show the diversity of the object classes involved in the attack and the randomness of the target categories.
Figure 11: The total of 10 randomly selected 3D adversarial objects under 10 random viewpoints, along with their predicted outcomes. This validates the randomness and diversity of our viewpoint selection, as well as the robustness of our TT3D method across varying viewpoints.
B.4 Transferability in the Physical World
Our TT3D could achieve a transferable targeted attack not only in the digital world but also in the physical world. The specific results in the digital world have been given in the manuscript. Here, to verify the performance in the physical world, we follow the implementation process in Appendix A and test our 3D-printed adversarial objects’ attack success rate against various black-box classifiers in the physical world. As shown in Tab. 7, even when confronted with varying backgrounds, our 3D adversarial objects still maintain commendable attack success rates across a spectrum of black-box classifiers, which verifies the robustness of our TT3D.
Appendix C More examples of TT3D
In this section, we present more 3D adversarial samples generated by TT3D: (1) Fig. 10 shows 100 different 3D adversarial objects, generated by attacking the surrogate model ResNet-101. These images, captured from a single random viewpoint, demonstrate the diversity of the object classes involved in the attack and the randomness of the target categories. (2) Fig. 11 consists of 10 randomly selected 3D objects from Fig. 10, each depicted from 10 different viewpoints (100 random viewpoints used in the experiments), to illustrate the effectiveness of the 3D adversarial samples under various viewpoints.